# On the Identifiability of Tensor Ranks via Prior Predictive Matching

### v1

@@ B.3.4 General Guideline
Appendix C Further Experiments
@@
Across settings, TT rank recovery tracks the y=xy{=}x line for small and moderate true ranks, with dispersion increasing as ranks grow or the signal-to-noise ratio (SNR) drops (e.g., smaller Poisson means or more dispersed Gamma priors). This is expected for moment-based estimators: the pure terms $v_{S}$ shrink in magnitude at low SNR, making inclusion–exclusion noisier and the denominators in the closed-form ratios closer to zero, which amplifies variance and occasionally induces slight upward bias. Interior bonds are typically better behaved than boundary bonds, which involve longer interaction sets. One practical mitigations help in our runsL mean normalization of all $\hat{v}_{S}$ by $(\widehat{E}[Y])^{2}$, which reduces spread without changing the target. Across panels, estimation accuracy improves as the effective SNR increases: low-mean or high-variance priors yield underestimation and wide dispersion, while moderate $(\mu,\mathrm{cv})$ combinations (as predicted by the SNR sweet-spot analysis in Appendix B.2) yield nearly unbiased recovery. The results confirm that the covariance-only TT estimator remains consistent when the prior mean and variance are chosen to keep $\mathrm{SNR}(S)\!\gtrsim\!5$ for the pure-term subsets involved in the ratio identities.

### v2

@@ B.3.4 General Guideline
Appendix C ADDITIONAL EXPERIMENTS
@@
Across settings, TT rank recovery tracks the y=xy{=}x line for small and moderate true ranks, with dispersion increasing as ranks grow or the signal-to-noise ratio (SNR) drops (e.g., smaller Poisson means or more dispersed Gamma priors). This is expected for moment-based estimators: the pure terms $v_{S}$ shrink in magnitude at low SNR, making inclusion–exclusion noisier and the denominators in the closed-form ratios closer to zero, which amplifies variance and occasionally induces slight upward bias. Interior bonds are typically better behaved than boundary bonds, which involve longer interaction sets. One practical mitigation that helped in our runs was mean normalization of all $\hat{v}_{S}$ by $(\widehat{E}[Y])^{2}$, which reduces spread without changing the target. Across panels, estimation accuracy improves as the effective SNR increases: low-mean or high-variance priors yield underestimation and wide dispersion, while moderate $(\mu,\mathrm{cv})$ combinations (as predicted by the SNR sweet-spot analysis in Appendix B.2) yield nearly unbiased recovery. The results confirm that the covariance-only TT estimator remains consistent when the prior mean and variance are chosen to keep $\mathrm{SNR}(S)\!\gtrsim\!5$ for the pure-term subsets involved in the ratio identities.